OpenStudy (anonymous):

Help!! Can someone help me solve. -5/2(1/4)+7/8/5/6

OpenStudy (texaschic101):

-5/2(1/4) + (7/8) / (5/6) - 5/8 + (7/8) / (5/6) -5/8 + 7/8 * 6/5 -5/8 + 42/40 -25/40 + 42/40 17/40 ==> I got this

OpenStudy (anonymous):

Ugh sigh, its saying it \[\frac{ 29 }{ 42 }\]

OpenStudy (texaschic101):

is the problem : -5 / (2 * 1/4) + (7/8) / (5/6) or (-5/2)(1/4) + (7/8) / (5/6)

OpenStudy (texaschic101):

@whpalmer4 .....I do not see what I did wrong ?

OpenStudy (whpalmer4):

uh, could you draw the problem, or use the equation editor button? I have no idea how to interpret this properly...

OpenStudy (anonymous):

I don't either :-(

OpenStudy (whpalmer4):

but you see it on the page in your book or problem set or screen, and I'm sure it doesn't look exactly like -5/2(1/4)+7/8/5/6 there.

OpenStudy (anonymous):

How about: \[-\frac{ 5 }{ 6 }\left(\begin{matrix}\frac{ 1 }{ 2 } \\ \end{matrix}\right)+\frac{ 7 }{ 9 }\div \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

Sorry I never used the equation options before

OpenStudy (anonymous):

I got this wrong as well I thought the answer was 67/72

OpenStudy (whpalmer4):

Okay, here is the problem that gives that result :-) \[-\frac{5}{6}\left(\frac{1}{4}\right) + \frac{7}{9}/\frac{5}{6}\]

OpenStudy (whpalmer4):

oh, wait, this gives 29/40, not 29/42, could you check to make sure of the correct answer?

OpenStudy (anonymous):

It says correct answer is \[-\frac{ 19 }{ 60 }\]

OpenStudy (anonymous):

shout I'm sorry guys the -5/6 should be -5/2

OpenStudy (whpalmer4):

to find the least common denominator, write each denominator as a product of primes: 12 = 2*2*3 = 2^2*3 45 = 3*3*5 = 3^2*5 the LCD will be the highest power of each term found: 2^2*3^2*5 = 180

OpenStudy (texaschic101):

-5/2(1/2) + (7/9) / (5/6) -5/4 + 7/9 * 6/5 -5/4 + 42/45 -225/180 + 168/180 - 57/180

OpenStudy (anonymous):

I'm trying to figure out how the correct answer is -19/60

OpenStudy (texaschic101):

-57/180 reduces to -19/60

OpenStudy (whpalmer4):

19*3 = 57

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Thank you guys for your help. I going to bed :-)

OpenStudy (texaschic101):

night - night :)